Please Answer the questions:

1. How is the future value of a single cash flow computed?

2. How is the present value of a series of cash flows computed.

3. What is the Net Present Value of an investment?

4. What is an EAR, and how is it computed?

5. What is a perpetuity? An annuity?
Answer #1

FV = Future Value

PV = Present Value

r = discount rate

n = number of periods or number of payments in case of annuity.

1.

FV = PV * (1+r)^{n}

2.

PV = \frac{PMT}{r} * \left [ 1 - \frac{1}{(1+r)^{n}} \right ]

3. Net Present Value of an investment is the present value of all the cash inflows and outflows of the investment. The present value is calculated by discounting the cash flows using a rate, that reflects the risk of the cash flows.

4. EAR is the Effective Annual Rate

It means that the interest rate is annual and the frequency of compounding is also annual.

Eg. 12% per year compounded annually.

EAR = ( 1 + r )^n - 1

Where r is another effective rate and n is the number of compoundings in a year.

5. If the cash flows last forever then the series of cash flows are called perpetuity Or an investment that produces a series of cash flows forever is called a perpetual investment.

On the other hand, if the cash flows are for only a specific period of time, then the investment is called annuity.

Can you please upvote? Thank You :-)